Find margin of error and confidence interval for the population mean µ of the college graduates who have taken the statistics course if average value for sample of 25 objects is 52,620 and sample standard deviation S=10,700 . Use 95% confidence interval. (Use t-distribution)



1
Expert's answer
2021-05-07T10:29:03-0400

Given, "n=25, \mu=56250,\sigma=10,700"


"\alpha=0.05, t_{\frac{\alpha}{2}}=t_{0.025}=2.064"


Critical value is 2.064.


Margin of Error (MOE)=critical value "\times" population standard deviation


"=2.064\times \dfrac{10700}{\sqrt{25}}"


"=2.064\times \dfrac{10700}{5}=4416.96"


Confidence Interval "=\mu\pm MOE=52620\pm 4416.96=(48233.04,57066.9)"